I would appreciate it if you can help me this is a Quadratic function question.
I would try to begin but I have no idea on how to start.


1) A ball is thrown upward with an initial velocity of 32 ft./sec from a height of six ft.
The function s(t)=-16t*+32t+6 gives the height of the ball t seconds after release. Determine the time when the ball is at maximum height, and find that height.

2) Jessica can be paid one of two ways for painting a house:
Plan A: $200 plus $12 per hour
Plan B: $20 per hour
Suppose the job takes n hours to complete. For what values of n is Plan A better for Jessica?

3) In a study, it was found that the threshold weight, W, defined as the weight above which the risk of death rises dramatically, is given by
W(h)=( h/12.3) cube
Find the threshold weight for someone 5 ft 7 inches tall.

Thank you for your help!

Answer by Alan3354(69443)

1) A ball is thrown upward with an initial velocity of 32 ft./sec from a height of six ft.
The function s(t)=-16t*+32t+6 gives the height of the ball t seconds after release. Determine the time when the ball is at maximum height, and find that height.
Method 1:
Set the 1st derivate = 0.
-32t + 32 = 0
t = 1 second
s(1) = -16 + 32 + 6
max = 22 feet
------------
Method 2:
pick a height, eg, the starting height of 6 feet and set it equal to that.
-16t^2 + 32t + 6 = 6
16t^2 = 32t
t = 0, t = 2
These are the 2 times that the ball is at 6 feet. The center time, 1 sec, is the time the ball is at its max.
(0+2)/2 = 1 second, same answer.
---------------------------
I'll do the others later.
------------------------
2) Jessica can be paid one of two ways for painting a house:
Plan A: $200 plus $12 per hour
Plan B: $20 per hour
Suppose the job takes n hours to complete. For what values of n is Plan A better for Jessica?

3) In a study, it was found that the threshold weight, W, defined as the weight above which the risk of death rises dramatically, is given by
W(h)=( h/12.3) cube
Find the threshold weight for someone 5 ft 7 inches tall.
Is h in inches?
W(67) = (67/12.3)^3
= 161.6 (is that pounds?)
You need to specify units.